circuit. VII Find the renin Thevenin equivalent 2v 6oon + 5 3002 3 Des158 @ q cos(500+) (2) I iso mF
Answer #1

Solution: Zth:- Thevenin Jurpedance, for calculating 2 th , we have to deactivate independent voltage sources, and dopen doutVe (1+2j) + 2V2 = 600 - 0 V₂ - V q = 24 . As Vevey VZ-V4 = 21 ) Vq = V2 Putting value of v= uz to in eo V₂ (1+28) + 2V 2 = 60Now, VT: Thevenir voltage 2V 6oor ve a Lo v % v izoor 3000 Vt 963 (500+) = 9 Lov using Nodal analysis at node - q LO + 600 +Then. -Vq - 2V TV The co The 2V TV, As V=V, Then Vths 3V, = 3x2.236 / -15.945º V z 3.7087 (-15.945º v THE VENINS EQo CRT. zt